# Position: LLMs Can be Good Tutors in Foreign Language Education

### v1

@@ Discussion.
While LLM-based data reformation can significantly enhance FLE, several gaps warrant attention. Most current studies prioritize textual forms or single-modal approaches, which may overlook valuable multimodal resources such as interactive video and audio-based content (Ghosal et al., 2023). Furthermore, cultural adaptation, although promising, remains underexplored in practical classroom scenarios, particularly for underrepresented personas and culturally sensitive topics. AlKhamissi et al. (2024) demonstrate how cultural misalignment can increase bias, suggesting Anthropological Prompting to strengthen cultural alignment. However, robust empirical evaluations are still limited across diverse learners and linguistic backgrounds. These issues highlight the importance of interdisciplinary research, where educators, linguists, and technologists collaborate to refine data reformation techniques that broaden inclusivity and effectiveness in FLE.
@@ 4.3 Data Annotation
While Data Creation focuses on generating learner-specific data, it often prioritizes diversity and adaptability over precision. The approach is particularly useful for tasks with large label spaces (Ding et al., 2024). In contrast, Data Annotation emphasizes producing high-quality, meticulously labeled data that is essential for tasks requiring accuracy and consistency. Unlike data creation, annotated data often undergoes rigorous validation to ensure its accuracy and relevancy (Artemova et al., 2024).
@@
LLMs can be central to generating a variety of annotations, including categorical labels, rationales, pedagogical feedback, and linguistic features such as discourse relations. Recent prompt engineering and fine-tuning techniques have further expanded LLMs’ annotation capabilities. For instance, Ye et al. (2024) leverage GPT-4 to annotate structured explanations for Chinese grammatical error correction, while Samuel et al. (2024) examine GPT-4 as a surrogate for human annotators in low-resource reading comprehension tasks. Likewise, Siyan et al. (2024) deploy GPT-4-Turbo for audio transcript annotations. However, inconsistencies across LLMs (Törnberg, 2024) remain a serious challenge, posing risks to educational reliability.
@@
Although LLM-based annotation generation is efficient, it also raises critical issues of bias, calibration, and validity—particularly in low-resource language contexts (Bhat & Varma, 2023; Jadhav et al., 2024). Automated or semi-automated evaluation strategies have emerged to address these quality concerns. For example, LLMs-as-Judges (Li et al., 2024a, b; Gu et al., 2024) reduce human overhead by automating evaluation, an approach increasingly explored in education-focused applications (Chiang et al., 2024; Zhou et al., 2024). However, purely automated frameworks can still propagate errors or bias.
@@ Discussion.
Although LLMs provide efficient data annotation, the inconsistency across different models remains a critical concern, affecting the quality and reliability of annotated educational materials. These discrepancies hinder the creation of universally reliable FLE content, especially in diverse linguistic and cultural contexts. Additionally, automated annotations often lack the nuance needed for pedagogical applications, making it essential to involve human oversight in critical cases to mitigate errors or biases.
Our position. We acknowledge the current limitations in LLM-based data creation, reformation, and annotation for FLE. However, we believe that with continued interdisciplinary collaboration, including input from educators, linguists, and technologists, these challenges can be addressed. Future advancements should focus on enhancing the accuracy and diversity of generated content, improving multimodal and culturally sensitive learning materials, and integrating more robust systems for human-LLM collaboration (Li et al., 2023; Wang et al., 2024e) in data annotation. This approach will ensure that LLMs can fully realize their potential as effective tutors in FLE.
@@ 5 LLMs as Task Predictors
Task-Based Language Learning (TBLL) (Nunan, 1989; Willis, 2021) as a methodological approach is one of the critical factors of modern foreign language teaching. LLMs have demonstrated remarkable capabilities in understanding and generating human language, making them well-suited for addressing numerous tasks in FLE. These tasks can be broadly categorized into three types based on their nature and the role of LLMs: 1) Discriminative, 2) Generative, and 3) Mixed of the above two roles.
@@ 5.1 Discriminative Task Predictors
Discriminative tasks in FLE primarily involve classifying learner inputs or grading their future performance. Below are some applications that are still calling for improvements:
@@
The task aims to automatically grade students’ assignments, including essay scoring (Seßler et al., 2024; Li & Liu, 2024; Syamkumar et al., 2024), short answer grading (Schneider et al., 2023; Henkel et al., 2024), and spoken language evaluation (Gao et al., 2023; Fu et al., 2024). LLMs can process learners’ submissions to judge grammar, lexical diversity, coherence, and even spoken fluency, providing instant feedback. This scalability is particularly appealing for large classes, where human evaluators are often overwhelmed and unable to provide timely, personalized critique (Mizumoto & Eguchi, 2023).
@@
Given sequences of learning interactions in online learning systems, Knowledge Tracing identifies and tracks students’ evolving mastery of target skills (Shen et al., 2024b; Xu et al., 2023). LLM-based methods have been explored in cold-start scenarios (Zhan et al., 2024; Jung et al., 2024), offering strong generalization by inferring latent learner states from limited data. These approaches can support adaptive learning pathways, giving personalized recommendations based on predicted performance and knowledge gaps.
@@ Discussion.
Despite their promise in automating and personalizing these discriminative tasks, LLMs still grapple with notable limitations that hinder their utility as robust tutoring tools. First, misalignment of assessment with expert instructors poses risks: machine-generated scores may deviate from established rubrics or neglect qualitative nuances, leading to potential discrepancies in grading quality (Kundu & Barbosa, 2024). Second, the lack of empathy compounds this issue, as assessments devoid of human judgment risk discouraging learners or overlooking subtle motivational factors (Sharma et al., 2024). Knowledge tracing approaches, while promising in cold-start scenarios, struggle with capturing the complexity of long-term learning trajectories and deeper cognitive processes (Cho et al., 2024). These concerns point to the need for more transparent and human-centered methods in utilizing LLMs for assessment.
@@ 5.2 Generative Task Predictors
Generative tasks involve producing new content or responses. LLMs are known to be adept in these tasks due to their natural language generation capabilities.
@@
In foreign language writing, errors often reveal learners’ gaps in grammar and vocabulary (Hyland & Hyland, 2006). LLMs can detect and correct these errors (Bryant et al., 2023; Ye et al., 2023), offering concise explanations (Ye et al., 2024) that reinforce language rules. By streamlining error detection and pedagogically framing corrections, learners deepen their linguistic understanding.
@@
Quizzes and exercises remain vital in FLE for practice and targeted remediation (Rashov, 2024). LLMs enhance this process by delivering prompt, personalized feedback that pinpoints strengths and addresses weaknesses (Borges et al., 2024). This scalability enables learners to self-regulate and refine their skills without relying solely on human graders (Stamper et al., 2024).
@@
Moving beyond straightforward Q&A, Socratic questioning promotes critical thinking and self-reflection (Paul & Elder, 2007). SocraticLM (Liu et al., 2024b), for example, aligns an LLM with open-ended, inquiry-based teaching principles, guiding learners through iterative exploration rather than prescriptive correction. In theory, this fosters deeper conceptual understanding and active learner engagement.
@@ Discussion.
Despite the promise of LLM-based generation in FLE, multiple uncertainties persist. Determining how to provide automatic feedback that genuinely maximizes learning outcomes is an ongoing challenge (Stamper et al., 2024), particularly given education’s risk-averse culture and high accountability standards (Xiao et al., 2024). Moreover, while LLMs like SocraticLM have demonstrated success in domains like mathematics, their applicability to FLE contexts has not been thoroughly validated (Liu et al., 2024b). As such, the design of strategies and follow-up queries remains an open question in ensuring that these systems track and respond to learners’ cognitive states effectively.
@@ 5.3 Mixed Task Predictors
While discriminative systems for essay scoring and speech evaluation primarily focus on assigning grades, LLMs extend these capabilities by simultaneously generating formative feedback (Katuka et al., 2024; Stahl et al., 2024b). For example, an LLM can evaluate the coherence and lexical diversity of a written assignment, then offer specific revision strategies. In speaking practice, it can measure fluency and pronunciation accuracy while suggesting drills to refine intonation or stress patterns. Through this combination of scoring and tailored advice, learners gain a deeper understanding of their strengths and areas for improvement.
@@
Error Analysis systematically uncovers and categorizes learners’ missteps, from syntactic lapses in writing to flawed pronunciations in speaking (James, 2013; Erdoğan, 2005). LLMs functioning in a mixed capacity can classify these errors and generate corrective guidance, providing revised sentences, clarifications of grammatical rules, or remediation exercises for identified weaknesses (Myles, 2002; Mashoor & Abdullah, 2020). Such insight facilitates targeted interventions that enhance language proficiency across modalities, including reading and listening.
@@ Discussion.
Mixed-task systems hold promise by combining assessment and feedback generation, but they face notable challenges. One major issue is the weak alignment between scoring mechanisms and the quality of feedback provided (Stahl et al., 2024b). For example, while essay scoring systems may deliver comprehensive evaluations, the feedback often lacks specificity, limiting its instructional value. Additionally, although error analysis has potential, the absence of standardized pedagogical benchmarks, especially in oral tasks, hampers the reliability and comparability of LLM-based tools (Leu Jr, 1982). Furthermore, by addressing errors across these modalities, LLMs can deliver a holistic and adaptive learning experience that fosters deeper language proficiency (Zhao et al., 2024).
Our position. While LLMs offer scalable solutions for task prediction in FLE, their current limitations—such as misalignment with expert assessments, lack of empathy, and weak alignment between assessment and feedback—require ongoing refinement. Future research should focus on improving model transparency, enhancing the cultural and emotional sensitivity of LLMs, and refining task predictors to better reflect long-term learning trajectories and learner motivation. Additionally, developing standardized pedagogical benchmarks for error analysis, especially in oral tasks, will help ensure the consistency and reliability of LLM-generated feedback.
@@ 6 LLM-empowered Agent
In this section, we delve into the potential of LLMs as intelligent agents in FLE. LLMs can act as catalysts for personalized learning, addressing the long-standing scalability, adaptability, and inclusivity challenges in traditional teaching paradigms.
@@ 6.1 Fundamental Abilities
LLMs excel at merging structured educational knowledge graphs (Abu-Rasheed et al., 2024; Hu & Wang, 2024) with unstructured textual data (Li et al., 2024c; Modran et al., 2024), providing rich, contextualized information on linguistic constructs and cultural nuances. Their ability to perform real-time knowledge editing (Wang et al., 2024d; Zhang et al., 2024a) ensures learners receive content aligned with evolving language usage, addressing the inherent limitations of static materials.
@@
LLMs require embedding with pedagogical principles to facilitate genuine learning experiences (Carroll, 1965; Taneja, 1995). Recent work incorporates theoretical frameworks, such as Bloom’s taxonomy (Bloom et al., 1956), to guide LLMs in systematically addressing different cognitive levels (Jiang et al., 2024b). Approaches like Pedagogical Chain of Thought (Jiang et al., 2024b) and preference learning (Sonkar et al., 2024; Rafailov et al., 2024) focus on aligning model responses with educational objectives.
@@
By assisting in crafting teaching objectives and lesson designs, LLMs can handle complex tasks such as differentiated instruction (Hu et al., 2024). LessonPlanner (Fan et al., 2024) has been proposed to assist novice teachers in preparing lesson plans, with expert interviews confirming its effectiveness. Zheng et al. (2024) propose a three-stage process to produce customized lesson plans, using Retrieval-Augmented Generation (RAG), self-critique, and subsequent refinement.
@@
Effective tutoring systems track learner histories and tailor subsequent interactions accordingly (Jiang et al., 2024a; Chen et al., 2024). When serving as memory-augmented agents, LLMs can retain individualized data—such as repeated grammar mistakes or overlooked vocabulary—thereby improving continuity and enabling consistent scaffolding of future learning tasks.
@@
Beyond textual interactions, LLM-based agents can integrate specialized tools to streamline the educational ecosystem, from cognitive diagnosis modules (Ma & Guo, 2019) to report generators (Zhou et al., 2025). By orchestrating these resources, LLMs seamlessly unify diverse utilities under a single interface, enhancing learner experience and instructional efficiency.
@@ Discussion.
Existing research often overlooks the interplay among listening, speaking, reading, and writing in real-world language learning (Caines et al., 2023; Shetye, 2024). Most systems concentrate on text-oriented features, lacking robust benchmarks and methodologies to evaluate integrated, multimodal interactions. Recent work on pedagogical alignment (Razafinirina et al., 2024) largely addresses textual data, leaving out real-time speaking and listening dynamics that demand complex, rapid feedback. Similarly, while memory modules can track repeated written errors, their effectiveness in monitoring and improving learning efficacy remains underexplored.
6.2 Applications
Although still in its early stages, LLM-empowered agents have already started to show promising applications in FLE.
Classroom Simulation.
Classroom simulation leverages LLM-empowered agents to recreate complex, interactive learning settings without the logistical hurdles of organizing physical classrooms (Zhang et al., 2024b). By simulating virtual students and tutors, researchers can study pedagogical strategies at scale, generate diverse learner interactions, and refine teaching techniques. Moreover, this virtual data can be used to fine-tune LLMs for specific educational contexts and learner profiles (Liu et al., 2024b), offering a cost-effective and adaptable approach to language instruction.
Intelligent Tutoring System (ITS).
LLM-based agents have demonstrated the capacity to provide dynamic, personalized tutoring experiences (Kwon et al., 2024), effectively identifying learner weaknesses through large-scale linguistic analysis (Caines et al., 2023). This makes them promising for delivering individualized instruction at scale. Although current ITS applications in mathematics (Pal Chowdhury et al., 2024) and science (Stamper et al., 2024) have shown success, the extension to FLE requires nuanced handling of cultural and contextual elements, as well as the unpredictability of human language usage.
Discussion.
Despite the promise of these applications, critical challenges remain. Existing classroom simulation frameworks often lack standardized benchmarks for FLE, making it difficult to assess the efficacy and generalizability of developed systems (Zhang et al., 2024b). In addition, evaluating language-specific tutoring strategies, including real-time conversational practice and holistic skill integration, remains an underexplored frontier. Addressing these gaps requires new datasets and metrics centered on holistic skill development, as well as interdisciplinary collaboration.
Our position. We argue that to overcome current limitations, future research should focus on the integration of multimodal learning tasks (Sonlu et al., 2024) and the development of standardized frameworks for evaluating FLE simulations. Moreover, LLMs should evolve beyond text-based capabilities to provide real-time, context-sensitive feedback, particularly in speaking and listening. Interdisciplinary collaboration and the creation of new datasets tailored to FLE are crucial for refining these systems and ensuring their scalability and inclusivity in language instruction. Additionally, addressing the complexities of cultural context and learner variability will be key to the success of LLMs as effective agents in FLE.
7 Challenges and Future Directions
While we posit that LLMs have the potential to revolutionize FLE, realizing their full promise requires addressing key challenges. This section offers a concise overview of these challenges, followed by directions that could guide future research and deployment.

### v2

@@ Discussion.
While LLM-based data reformation can significantly enhance English Education, several gaps warrant attention. Most current studies prioritize textual forms or single-modal approaches, which may overlook valuable multimodal resources such as interactive video and audio-based content Ghosal et al. (2023). Furthermore, cultural adaptation, although promising, remains underexplored in practical classroom scenarios, particularly for underrepresented persons and culturally sensitive topics. AlKhamissi et al. (2024) demonstrate how cultural misalignment can increase bias. However, robust empirical evaluations are still limited across diverse learners and linguistic backgrounds.
@@ 4.3 Data Annotation
While Data Creation focuses on generating learner-specific data, it often prioritizes diversity and adaptability over precision. The approach is particularly useful for tasks with large label spaces Ding et al. (2024). In contrast, Data Annotation emphasizes producing high-quality, meticulously labeled data that is essential for tasks requiring accuracy and consistency. Unlike data creation, annotated data often undergoes rigorous validation to ensure its accuracy and relevancy Artemova et al. (2024).
@@
LLMs can be central to generating a variety of annotations, including categorical labels, rationales, pedagogical feedback, and linguistic features such as discourse relations. Recent prompt engineering and fine-tuning techniques have further expanded LLMs’ annotation capabilities. For instance, Ye et al. (2024) leverage GPT-4 to annotate structured explanations for Chinese grammatical error correction, while Samuel et al. (2024) examine GPT-4 as a surrogate for human annotators in low-resource reading comprehension tasks. Likewise, Siyan et al. (2024) deploy GPT-4-Turbo for audio transcript annotations. However, inconsistencies across LLMs Törnberg (2024) remain a serious challenge, posing risks to educational reliability.
@@
Although LLM-based annotation is efficient, it also raises critical issues of bias, calibration, and validity, particularly in low-resource language contexts Bhat and Varma (2023); Jadhav et al. (2024). Automated or semi-automated evaluation strategies have emerged to address these quality concerns. For example, LLMs-as-Judges Li et al. (2024a, b); Gu et al. (2024) reduce human overhead by automating evaluation, an approach increasingly explored in education-focused applications Chiang et al. (2024); Zhou et al. (2024). However, purely automated frameworks can still propagate errors or bias.
@@ Discussion.
Although LLMs provide efficient data annotation, the inconsistency across different models remains a critical concern, affecting the quality and reliability of annotated educational materials. These discrepancies hinder the creation of universally reliable educational content, especially in diverse linguistic and cultural contexts. Additionally, automated annotations often lack the nuance needed for pedagogical applications, making it essential to involve human oversight in critical cases to mitigate errors or biases.
Our position. We acknowledge the current limitations in LLM-based data creation, reformation, and annotation for English Education. However, we believe that with continued interdisciplinary collaboration, these challenges can be addressed. Future advancements should focus on enhancing the accuracy and diversity of generated content, improving multi-modal and culturally sensitive learning materials, and integrating more robust systems for human-LLM collaboration Li et al. (2023); Wang et al. (2024e) in data annotation. This will ensure that LLMs can fully realize their potential as effective tutors in English Education.
@@ 5 LLMs as Task Predictors
Task-Based Language Learning (TBLL) Nunan (1989); Willis (2021) as a methodological approach is one of the effective English Education methods. LLMs have demonstrated remarkable capabilities in understanding and generating human language, making them well-suited for addressing numerous tasks in English Education. These tasks can be broadly categorized into three types based on their nature and the role of LLMs: 1) Discriminative, 2) Generative, and 3) Mixed of the above two roles.
@@ 5.1 Discriminative Task Predictors
Discriminative tasks in English Education primarily involve classifying learner inputs or grading their future performance. Below are some applications that are still calling for improvements:
@@
The task aims to automatically grade students’ assignments, including essay scoring Seßler et al. (2024); Li and Liu (2024); Syamkumar et al. (2024), short answer grading Schneider et al. (2023); Henkel et al. (2024), and spoken language evaluation Gao et al. (2023); Fu et al. (2024). LLMs can process learners’ submissions to judge grammar, lexical diversity, coherence, and even spoken fluency, providing instant feedback. This scalability is particularly appealing for large classes, where human evaluators are often overwhelmed and unable to provide timely, personalized critique Mizumoto and Eguchi (2023).
@@
Given sequences of learning interactions in online learning systems, Knowledge Tracing identifies and tracks students’ evolving mastery of target skills Shen et al. (2024b); Xu et al. (2023). LLM-based methods of Knowledge Tracing have been explored in cold-start scenarios Zhan et al. (2024); Jung et al. (2024), offering strong generalization by inferring latent learner states from limited data. These approaches can support adaptive learning pathways, giving personalized recommendations based on predicted performance and knowledge gaps.
@@ Discussion.
Despite their promise in automating and personalizing these discriminative tasks, LLMs still grapple with notable limitations that hinder their utility as robust tutoring tools. First, misalignment of assessment with expert instructors poses risks: machine-generated scores may deviate from established rubrics or neglect qualitative nuances, leading to potential discrepancies in grading quality Kundu and Barbosa (2024). Second, the lack of empathy compounds this issue, as assessments devoid of human judgment risk discouraging learners or overlooking subtle motivational factors Sharma et al. (2024). Knowledge tracing approaches, while promising in cold-start scenarios, struggle with capturing the complexity of long-term learning trajectories and deeper cognitive processes Cho et al. (2024). These concerns point to the need for more transparent and human-centered methods in utilizing LLMs for assessment.
@@ 5.2 Generative Task Predictors
Generative tasks involve producing new content or responses. LLMs are known to be adept at these tasks due to their generation capabilities.
@@
In English writing, errors often reveal learners’ gaps in grammar and vocabulary Hyland and Hyland (2006). LLMs can detect and correct these errors Bryant et al. (2023); Ye et al. (2023), offering concise explanations Ye et al. (2024) that reinforce language rules. By streamlining error detection and corrections, learners deepen their linguistic understanding.
@@
Quizzes and exercises remain vital in English Education for practice and targeted remediation Rashov (2024). LLMs enhance this process by delivering prompt, personalized feedback that pinpoints strengths and addresses weaknesses Borges et al. (2024). This scalability enables learners to self-regulate and refine their skills without relying solely on human graders Stamper et al. (2024).
@@
Moving beyond straightforward Q&A, Socratic questioning promotes critical thinking and self-reflection Paul and Elder (2007). SocraticLM Liu et al. (2024b), for example, aligns an LLM with open-ended, inquiry-based teaching principles, guiding learners through iterative exploration rather than prescriptive correction. In theory, this fosters deeper conceptual understanding and active learner engagement.
@@ Discussion.
Despite the promise of LLM-based generation in English Education, multiple uncertainties persist. Determining how to provide automatic feedback that genuinely maximizes learning outcomes is an ongoing challenge Stamper et al. (2024), particularly given education’s risk-averse culture and high accountability standards Xiao et al. (2024). Moreover, while LLMs like SocraticLM have demonstrated success in domains like mathematics, their applicability to English Education contexts has not been thoroughly validated Liu et al. (2024b). As such, the design of strategies and follow-up queries remains an open question in ensuring these systems track and respond to learners’ cognitive states.
@@ 5.3 Mixed Task Predictors
While discriminative systems for automated essay scoring and speech evaluation primarily focus on assigning grades, LLMs extend these capabilities by simultaneously generating formative feedback Katuka et al. (2024); Stahl et al. (2024b). For example, an LLM can evaluate the coherence and lexical diversity of a written essay, then offer specific revision strategies. In speaking practice, it can measure fluency and pronunciation accuracy while suggesting drills to refine intonation or stress patterns. Through this combination of scoring and tailored advice, learners gain a deeper understanding of their strengths and areas for improvement.
@@
Error Analysis systematically uncovers and categorizes learners’ missteps, from syntactic lapses in writing to flawed pronunciations in speaking James (2013); Erdoğan (2005). LLMs functioning in a mixed capacity can classify these errors and generate corrective guidance, providing revised sentences, clarifications of grammatical rules, or remediation exercises for identified weaknesses Myles (2002); Mashoor and Abdullah (2020). Such insight facilitates targeted interventions that enhance language proficiency across modalities, including reading and listening.
@@ Discussion.
Mixed-task systems hold promise by combining assessment and feedback generation, but they face notable challenges. One major issue is the weak alignment between scoring mechanisms and the quality of feedback provided Stahl et al. (2024b). For example, while essay scoring systems may deliver comprehensive evaluations, the feedback often lacks specificity, limiting its instructional value. Additionally, although error analysis has potential, the absence of standardized pedagogical benchmarks, especially in oral tasks, hampers the reliability and comparability of LLM-based tools Leu Jr (1982).
Our position. While LLMs offer scalable solutions for task prediction in English Education, their current limitations—such as misalignment with expert assessments, lack of empathy, and weak alignment between assessment and feedback—require ongoing refinement. Future research should focus on improving model transparency, enhancing the cultural and emotional sensitivity of LLMs, and refining task predictors to better reflect long-term learning trajectories and learner motivation. Additionally, developing standardized pedagogical benchmarks for error analysis will help ensure the consistency and reliability of LLM-generated feedback.
@@ 6 LLM-empowered Agent
In this section, we delve into the potential of LLMs as intelligent tutoring agents in English Education. LLMs can act as catalysts for personalized learning, addressing the long-standing scalability, adaptability, and inclusivity challenges in traditional teaching paradigms.
@@ 6.1 Fundamental Abilities
LLMs excel at merging structured educational knowledge graphs Abu-Rasheed et al. (2024); Hu and Wang (2024) with unstructured textual data Li et al. (2024c); Modran et al. (2024), providing rich, contextualized information on linguistic constructs and cultural nuances. Their ability to perform real-time knowledge editing Wang et al. (2024d); Zhang et al. (2024a) ensures learners receive content aligned with evolving language usage, addressing the inherent limitations of static materials.
@@
LLMs require embedding with pedagogical principles to facilitate genuine learning experiences Carroll (1965); Taneja (1995). Recent work incorporates theoretical frameworks, such as Bloom’s taxonomy Bloom et al. (1956), to guide LLMs in systematically addressing different cognitive levels Jiang et al. (2024b). Approaches like Pedagogical Chain of Thought Jiang et al. (2024b) and preference learning Sonkar et al. (2024); Rafailov et al. (2024) focus on aligning model responses with educational objectives.
@@
By assisting in crafting teaching objectives and lesson designs, LLMs can handle complex tasks such as differentiated instruction Hu et al. (2024). LessonPlanner Fan et al. (2024) has been proposed to assist novice teachers in preparing lesson plans, with expert interviews confirming its effectiveness. Zheng et al. (2024) propose a three-stage process to produce customized lesson plans, using Retrieval-Augmented Generation (RAG), self-critique, and subsequent refinement.
@@
Effective tutoring systems track learner histories and tailor subsequent interactions accordingly Jiang et al. (2024a); Chen et al. (2024). When serving as memory-augmented agents, LLMs can retain individualized data—such as repeated grammar mistakes or overlooked vocabulary—thereby improving continuity and enabling consistent scaffolding of future learning tasks.
@@
Beyond textual interactions, LLM-based agents can integrate specialized tools to streamline the educational ecosystem, from cognitive diagnosis modules Ma and Guo (2019) to report generators Zhou et al. (2025). By orchestrating these resources, LLMs seamlessly unify diverse utilities under a single interface, enhancing learner experience and instructional efficiency.
6.2 Applications
Although still in its early stages, LLM-empowered agents have already started to show promising applications in English Education.
Classroom Simulation.
Classroom simulation leverages LLM-empowered agents to recreate complex, interactive learning settings without the logistical hurdles of organizing physical classrooms Zhang et al. (2024b). By simulating virtual students and tutors, researchers can study pedagogical strategies at scale, generate diverse learner interactions, and refine teaching techniques. Moreover, this virtual data can be used to fine-tune LLMs for specific educational contexts and learner profiles Liu et al. (2024b), offering a cost-effective and adaptable approach to language instruction.
Intelligent Tutoring System (ITS).
LLM-based agents have demonstrated the capacity to provide dynamic, personalized tutoring experiences Wang et al. (2025); Kwon et al. (2024), effectively identifying learner weaknesses through large-scale linguistic analysis Caines et al. (2023). This makes them promising for delivering individualized instruction at scale. Although current ITS applications in mathematics Pal Chowdhury et al. (2024) and science Stamper et al. (2024) have shown success, the extension to English Education requires nuanced handling of cultural and contextual elements, as well as the unpredictability of human language usage.
@@ Discussion.
Despite the promise of these applications, critical challenges remain. Existing classroom simulation frameworks often lack standardized benchmarks for English Education, making it difficult to assess the efficacy and generalizability of developed systems Zhang et al. (2024b). In addition, evaluating language-specific tutoring strategies, including real-time conversational practice and holistic skill integration, remains an underexplored frontier. Addressing these gaps requires new datasets and metrics centered on holistic skill development and interdisciplinary collaboration.
Our position. We argue that future research should focus on integrating multimodal learning tasks Sonlu et al. (2024) and developing standardized frameworks for evaluating English Education simulations. Moreover, LLMs should evolve beyond text-based capabilities to provide real-time, context-sensitive feedback, particularly in speaking and listening. Interdisciplinary collaboration and the creation of new datasets tailored to English Education are crucial for refining these systems and ensuring their scalability and inclusivity in language instruction.
7 Challenges
While we posit that LLMs have the potential to revolutionize English Education, realizing their full promise requires addressing key challenges. This section offers a concise overview of these challenges, followed by directions that could guide future research and deployment.